# EasyInv: Toward Fast and Better DDIM Inversion

### v1

@@ Abstract
This paper introduces EasyInv, an easy yet novel approach that significantly advances the field of DDIM Inversion by addressing the inherent inefficiencies and performance limitations of traditional iterative optimization methods. At the core of our EasyInv is a refined strategy for approximating inversion noise, which is pivotal for enhancing the accuracy and reliability of the inversion process. By prioritizing the initial latent state, which encapsulates rich information about the original images, EasyInv steers clear of the iterative refinement of noise items. Instead, we introduce a methodical aggregation of the latent state from the preceding time step with the current state, effectively increasing the influence of the initial latent state and mitigating the impact of noise. We illustrate that EasyInv is capable of delivering results that are either on par with or exceed those of the conventional DDIM Inversion approach, especially under conditions where the model’s precision is limited or computational resources are scarce. Concurrently, our EasyInv offers an approximate threefold enhancement regarding inference efficiency over off-the-shelf iterative optimization techniques.
Introduction
Diffusion models have become a major focus of research in recent years, mostly renowned for their ability to generate high-quality images that closely match given prompts. Among the many diffusion models introduced in the community, Stable Diffusion (SD) (Rombach et al. 2022) stands out as one of the most widely utilized in scientific research, largely due to its open-source nature. Another contemporary diffusion model gaining popularity is DALL-E 3 (Betker et al. 2023), which offers users access to its API and the ability to interact with it through platforms like ChatGPT (Openai 2024). These models have significantly transformed the visual arts industry and have attracted substantial attention from the research community.
While renowned generative diffusion models have made significant strides, a prevalent limitation is their reliance on textual prompts for input. This approach becomes restrictive when users seek to iteratively refine an image, as the sole reliance on prompts hinders flexibility. Although solutions such as ObjectAdd (Zhang, Lin, and Ji 2024) and P2P (Hertz et al. 2022) have been proposed to address image editing challenges, they are still confined to the realm of prompted image manipulation. Given that diffusion models generate images from noise inputs, a potential breakthrough lies in identifying the corresponding noise for any given image. This would enable the diffusion model to initiate the generation process from a known starting point, thereby allowing for precise control over the final output. The recent innovation of DDIM Inversion (Couairon et al. 2023) aims to overcome this challenge by reversing the denoising process to introduce noise. This technique effectively retrieves the initial noise configuration after a series of reference steps, thereby preserving the integrity of the original image while affording the user the ability to manipulate the output by adjusting the denoising parameters. With DDIM inversion, the generative process becomes more adaptable, facilitating the creation and subsequent editing of images with greater precision and control. For example, the MasaCtrl method (Cao et al. 2023) first transforms a real image into a noise representation and then identifies the arrangement of objects during the denoising phase. Portrait Diffusion(Liu et al. 2023) simultaneously inverts both the source and target images. Subsequently, it merges their respective $Q$, $K$ and $V$ values for mixups.
@@ Introduction
Figure 1: Performance comparison of inversion methods including vanilla DDIM Inversion (Couairon et al. 2023), Fixed-Point Iteration (Pan et al. 2023), ReNoise (Garibi et al. 2024) and our EasyInv. The proposed EasyInv performs well upon different diffusion models of SD-V1-4 and SD-XL.
@@
Considering the reliance on inversion techniques to preserve the integrity of the input image, the quality of the inversion process is paramount, as it profoundly influences subsequent tasks. As depicted in Figure 1(a), the performance of DDIM Inversion has been found to be less than satisfactory due to the discrepancy between the noise estimated during the inversion process and the noise expected in the sampling process. Consequently, numerous studies have been conducted to enhance its efficacy. In Null-Text inversion (Mokady et al. 2023), researchers observed that using a null prompt as input, the diffusion model could generate optimal results during inversion, suggesting that improvements to inversion might be better achieved in the reconstruction branch. Ju et al.’s work (Ju et al. 2023) exemplifies this approach by calculating the distance between latents at the current step and the previous step. PTI (Dong et al. 2023) opts to update the conditional vector in each step to guide the reconstruction branch for improving consistency. ReNoise (Garibi et al. 2024) focuses on refining the inversion process itself. This method iteratively adds and then denoises noise at each time step, using the denoised noise as input for the subsequent iteration. However, as shown in Figure 1(b), it can result in a black image output when dealing with certain special inputs, which will be discussed in detail in Sec. Experimentation. Pan et al. (Pan et al. 2023), while maintaining the iterative updating process, also amalgamated noise from previous steps with the current step’s noise. However, this method’s performance is limited in less effective models as displayed in Figure1(c). For instance, it performs well in SD-XL (Podell et al. 2023) but fails to yield proper results in SD-V1-4 (Rombach et al. 2022). We attribute this to their method’s sole focus on optimizing noise; when the noise is highly inaccurate, such simple optimization strategies encounter difficulties. Additionally, the iterative updating of noise is time-consuming, as Pan et al.’s method requires multiple model inferences per time step.
@@
In this paper, we conduct an in-depth analysis and recognize that the foundation of any inversion process is the initial latent state derived from a real image. Errors introduced at each step of the inversion process can accumulate, leading to a suboptimal reconstruction. Current methodologies, which focus on optimizing the transition between successive steps, may not be adequate to address this issue holistically. To tackle this, we propose a novel approach that considers the inversion process as a whole, underscoring the significance of the initial latent state throughout the process. Our approach, named EasyInv, incorporates a straightforward mechanism to periodically reinforce the influence of the initial latent state during the inversion. This is realized by blending the current latent state with the previous one at strategically selected intervals, thereby increasing the weight of the initial latent state and diminishing the noise’s impact. As a result, EasyInv ensures a reconstructed version that remains closer to the original image, as illustrated in Figure 1(d). Furthermore, by building upon the traditional DDIM Inversion framework (Couairon et al. 2023), EasyInv does not depend on iterative optimization between adjacent steps, thus enhancing computational efficiency. In Figure 2, we present a visualization of the latent states at the midpoint of the total denoising steps for various inversion methods. It is evident that the outcomes of our EasyInv are more closely aligned with the original image compared to all other methods, demonstrating that EasyInv achieves faster convergence.
@@
Figure 2: Visualization of the latent states midway through all denoising steps for various inversion methods. Our EasyInv shows its enhanced convergence by closely approximating the original image.
Related Works
@@
Diffusion Model. In recent years, there has been significant progress in the field of generative models, with diffusion models emerging as a particularly popular approach. The seminal denoising diffusion probabilistic models (DDPM) (Ho, Jain, and Abbeel 2020) introduced a practical framework for image generation based on the diffusion process. This method stands out from its predecessors, such as generative adversarial networks (GANs), due to its iterative nature. During the data preparation phase, Gaussian noise is incrementally added to a real image until it transitions into a state that is indistinguishable from raw Gaussian noise. Subsequently, a model can be trained to predict the noise added at each step, enabling users to input any Gaussian noise and obtain a high-quality image as a result. Ho et al(Ho, Jain, and Abbeel 2020) provided a robust theoretical foundation for their model, which has facilitated further advancements. Generative process in DDPM is both time-consuming and inherently stochastic due to the random noise introduced at each step. To address these limitations, the denoising diffusion implicit models (DDIM) were developed (Song, Meng, and Ermon 2020). By reformulating DDPM, DDIM has successfully reduced the amount of random noise added at each step. This reformulation results in a more deterministic denoising process. Furthermore, the absence of random noise allows for the aggregation of several denoising steps, thereby significantly reducing the overall computation time required to generate an image.
@@
Image Inversion. Converting a real image into noise is a pivotal first step in the realm of real image editing using diffusion models. The precision of this process has a profound impact in the final edit, with the critical element being the accurate identification of the noise added at each step. Couairon et al(Couairon et al. 2023) ingeniously swapped the roles of independent and implicit variables within the denoising function of the DDIM model, enabling it to predict the noise that should be introduced to the current latents. However, it is essential to recognize that the denoising step in a diffusion model is inherently an approximation, and when this approximation is utilized inversely, discrepancies between the model’s output and the actual noise value are likely to be exacerbated. To address this issue, ReNoise (Garibi et al. 2024) iterates through each noising step multiple times. For each inversion step, they employ an iterative approach to add and subsequently reduce noise, with the noise reduced in the final iteration being carried forward to the subsequent iteration. Pan et al. (Pan et al. 2023) offered a theoretical underpinning to the ReNoise method. Iterative optimization from ReNoise is classified under the umbrella of fixed-point iteration methods. Building upon Anderson’s seminal work (Anderson 1965), Pan et al. have advanced the field by proposing their novel method for optimizing noise during the inversion process.
Methodology
Preliminaries
DDIM Inversion
@@
Let $\mathbf{z}_{T}$ denote a noise tensor with $\mathbf{z}_{T}\sim\mathcal{I}(0,\mathbf{I})$. The DDIM (Couairon et al. 2023) leverages a pre-trained neural network $\varepsilon_{\theta}$ to perform $T$ denoising diffusion steps. Each step aims to estimate the underlying noise and subsequently restore a less noisy version of the tensor, $\mathbf{z}_{t-1}$, from its noisy counterpart $\mathbf{z}_{t}$ as:
@@
$\mathbf{z}_{t-1}=\sqrt{\frac{\alpha_{t-1}}{\alpha_{t}}}\mathbf{z}_{t}+\Bigg{(}\sqrt{\frac{1}{\alpha_{t-1}}-1}-\sqrt{\frac{1}{\alpha_{t}}-1}\Bigg{)}\cdot\varepsilon_{\theta}\big{(}\mathbf{z}_{t},t,\tau_{\theta}(y)\big{)},$ (1)
@@
where $t=T\rightarrow 1$, and $\{\alpha_{t}\}_{t=1}^{T}$ constitutes a prescribed variances set that guides the diffusion process. Furthermore, $\tau_{\theta}$ serves as an intermediate representation that encapsulates the textual condition $y$. For the convenience of following sections, we denote:
@@
Re-evaluating Eq. (1), we derive DDIM Inversion process (Couairon et al. 2023) as presented in Eq.(3). In this reformulation, we relocate an approximate $\mathbf{z}^{*}_{t}$ to the left-hand side, resulting in the following expression:
@@
Review. Given an image $\mathbf{I}^{*}$, after encoding it into the latent $\mathbf{z}^{*}_{0}$, we initiate $T$ inversion steps using Eq. (3) to obtain the noise $\mathbf{z}^{*}_{T}$. Starting with $\mathbf{z}_{T}=\mathbf{z}^{*}_{T}$, we proceed with a denoising process in Eq. (1) to infer an approximate reconstruction $\mathbf{z}_{0}$ that resembles the original latent $\mathbf{z}^{*}_{0}$. The primary source of error in this reconstruction arises from the difference between the noise predicted during the inversion process $\varepsilon_{\theta}\big{(}\mathbf{z}^{*}_{t-1},t-1,\tau_{\theta}(y)\big{)}$ and the noise expected in the sampling process, $\varepsilon_{\theta}\big{(}\mathbf{z}_{t},t,\tau_{\theta}(y)\big{)}$, denoted as $\varepsilon_{t}$, at each iterative step. This discrepancy originates from an imprecise approximation of the time step from $t$ to $t-1$. Therefore, reducing the discrepancy between the predicted noises at each step is crucial for achieving an accurate reconstruction, which is essential for the success of subsequent image editing tasks. For simplicity in the following expressions, we define:
$\varepsilon_{t}^{*}=\varepsilon_{\theta}\big{(}\mathbf{z}^{*}_{t-1},t-1,\tau_{\theta}(y)\big{)},\qquad\varepsilon_{t}=\varepsilon_{\theta}\big{(}\mathbf{z}_{t},t,\tau_{\theta}(y)\big{)}.$ (4)
Fixed-Point Iteration
@@
For clarity, let’s first restate Eq. (3) as follows:
@@
which represents the introduction of adding noise to the latent state $\mathbf{z}^{*}_{t-1}$. Under the assumption of Eq. (5), it should be the case that:
@@
Subsequently, by employing the noise estimation function from Eq. (2), we obtain:
@@
Given that $d(\mathbf{z}_{t})=\varepsilon_{t}$ and considering Eq. (5), we can deduce that:
@@
This formulation presents a fixed-point problem, which pertains to a value that remains unchanged under a specific transformation (Bauschke et al. 2011). In the context of functions, a fixed point is an element that is invariant under the application of the function. In this paper, we seek a $\varepsilon^{*}_{t}$ that, when transformed by $g$ and followed by $d$, can map back to itself, signifying an optimal solution as per Eq. (5).
@@
where $n$ denotes the iteration count. This iterative process can be enhanced through acceleration techniques such as Anderson acceleration  (Anderson 1965). However, calculating a complex $\varepsilon^{*}_{t}$ can be quite onerous. An empirical acceleration method proposed (Pan et al. 2023) introduces a refinement for $\varepsilon_{t}^{*}$ by setting: $\big{(}\varepsilon^{*}_{t})^{n}=\varepsilon_{\theta}(({\mathbf{z}^{*}_{t}})^{n},t-1,\tau_{\theta}(y)\big{)}$ and $(\varepsilon^{*}_{t})^{n-1}=\varepsilon_{\theta}\big{(}({\mathbf{z}^{*}_{t}})^{n-1},t-1,\tau_{\theta}(y)\big{)}$. They finally reach:
@@
where the term $0.5\cdot(\varepsilon^{*}_{t})^{n}+0.5\cdot(\varepsilon^{*}_{t})^{n-1}$ represents the refinement technique for $\varepsilon^{*}_{t}$ as suggested by Pan et al.. If we were to apply the function $d$ to both sides of Eq. (11), it would align perfectly with the form of Eq. (10). Their experiments have demonstrated that this approach is more effective than both Anderson’s method  (Anderson 1965) and other techniques in inversion tasks.
Despite the progress made, this paper acknowledges inherent limitations in the practical implementation of the inversion technique: (1) Inversion Efficiency: While the method outlined in Eq. (11) has shown improvements over traditional fixed-point iteration, it still relies on iterative optimization. The need for multiple forward passes through the diffusion model is computationally demanding and can result in inefficiencies in downstream applications. (2) Inversion Performance: The theoretical improvements presented assume that $\varepsilon^{*}_{t}=\varepsilon_{t}$. However, iterative optimization does not guarantee the exact fulfillment of Equation (7) for every time step $t$. Therefore, while the method may theoretically offer superior performance, cumulative errors can sometimes lead to practical outcomes that are less satisfactory than those achieved with the standard DDIM Inversion method, as shown in Figure 1.
EasyInv
To facilitate our subsequent analysis, we introduce the notation $\bar{\alpha}_{t}$ to represent $\sqrt{\frac{\alpha_{t}}{\alpha_{t-1}}}$ and $\bar{\beta}_{t}$ to denote $\sqrt{\frac{\alpha_{t}}{\alpha_{t-1}}}\Big{(}\sqrt{\frac{1}{\alpha_{t-1}}-1}-\sqrt{\frac{1}{\alpha_{t}}-1}\Big{)}$. With these notations, we can reframe Eq. (3) as follow:
$\mathbf{z}^{*}_{t}=\bar{\alpha}_{t}\mathbf{z}^{*}_{t-1}+\bar{\beta}_{t}\varepsilon^{*}_{t},$ (12)

### v4

@@ Abstract
This paper introduces EasyInv, an easy yet novel approach that significantly advances the field of DDIM Inversion by addressing the inherent inefficiencies and performance limitations of traditional iterative optimization methods. At the core of our EasyInv is a refined strategy for approximating inversion noise, which is pivotal for enhancing the accuracy and reliability of the inversion process. By prioritizing the initial latent state, which encapsulates rich information about the original images, EasyInv steers clear of the iterative refinement of noise items. Instead, we introduce a methodical aggregation of the latent state from the preceding time step with the current state, effectively increasing the influence of the initial latent state and mitigating the impact of noise. We illustrate that EasyInv is capable of delivering results that are either on par with or exceed those of the conventional DDIM Inversion approach, especially under conditions where the model’s precision is limited or computational resources are scarce. Concurrently, our EasyInv offers an approximate threefold enhancement regarding inference efficiency over off-the-shelf iterative optimization techniques. See code at https://github.com/potato-kitty/EasyInv.
Machine Learning, Diffusion Model, Inversion, ICML
@@
Figure 1: Comparison of inversion methods including vanilla DDIM Inversion (Couairon et al., 2023), Fixed-Point Iteration (Pan et al., 2023), ReNoise (Garibi et al., 2024) and our EasyInv.
1 Introduction
Diffusion models are renowned for their ability to generate high-quality images that closely match given prompts. Among the many diffusion models, Stable Diffusion (SD) (Rombach et al., 2022) stands out as one of the most widely utilized models. Another contemporary diffusion model gaining popularity is DALL-E 3 (Betker et al., 2023), which offers users access to its API and the ability to interact with it through platforms like ChatGPT (Openai, 2024). These models have transformed the visual arts industry and have attracted substantial attention from the research community. While renowned generative diffusion models have made strides, a prevalent limitation is their reliance on textual prompts for input. This approach becomes restrictive when users seek to iteratively refine an image, as the sole reliance on prompts hinders flexibility. Although solutions such as ObjectAdd (Zhang et al., 2024) and P2P (Hertz et al., 2022) have been proposed to address image editing challenges, they are still confined to the realm of prompted image manipulation. Given that diffusion models generate images from noise inputs, a potential breakthrough lies in identifying the corresponding noise for any given image. This would enable the diffusion model to initiate the generation process from a known starting point, thereby allowing for precise control over the final output. The recent innovation of DDIM Inversion (Couairon et al., 2023) overcomes this by reversing the denoising process to introduce noise. This technique retrieves the initial noise configuration after a series of reference steps, thereby preserving the integrity of the original image while affording the user the ability to manipulate the output by adjusting the denoising parameters. With DDIM inversion, the generative process becomes more adaptable, facilitating the creation and subsequent editing of images with greater precision and control. For example, MasaCtrl (Cao et al., 2023) first transforms a real image into a noise representation and then identifies the arrangement of objects during the denoising phase. Portrait Diffusion (Liu et al., 2023) inverts both the source and target images. It merges their respective $Q$, $K$ and $V$ values for mixups.
@@ 1 Introduction
Considering the reliance on inversion techniques to preserve the integrity of the input image, the quality of the inversion process is paramount, as it profoundly influences subsequent tasks. As depicted in Figure 1(a), the performance of DDIM Inversion has been found to be less than satisfactory due to the discrepancy between the noise estimated during the inversion process and the noise expected in the sampling process. Consequently, numerous studies have been conducted to enhance its efficacy. In Null-Text inversion (Mokady et al., 2023), researchers observed that using a null prompt as input, the diffusion model could generate optimal results during inversion, suggesting that improvements to inversion might be better achieved in the reconstruction branch. Ju et al.’s work (Ju et al., 2023) exemplifies this approach by calculating the distance between latents at the current step and the previous step. PTI (Dong et al., 2023) opts to update the conditional vector in each step to guide the reconstruction branch for improving consistency. ReNoise (Garibi et al., 2024) focuses on refining the inversion process itself. This method iteratively adds and then denoises noise at each time step, using the denoised noise as input for the subsequent iteration. However, as shown in Figure 1(b), it can result in a black image output when dealing with certain special inputs, which will be discussed in detail in Section 4. Pan et al. (Pan et al., 2023), while maintaining the iterative updating process, also amalgamated noise from previous steps with the current step’s noise. However, this method’s performance is limited in less effective models as displayed in Figure 1(c). For instance, it performs well in SD-XL (Podell et al., 2023) but fails to yield proper results in SD-V1-4 (Rombach et al., 2022). We attribute this to their sole focus on optimizing noise; when the noise is highly inaccurate, such simple optimization strategies encounter difficulties. Additionally, the iterative updating of noise is time-consuming, as Pan et al.’s method requires multiple model inferences per time step.
@@
In this paper, we conduct an in-depth analysis and recognize that the foundation of any inversion process is the initial latent state derived from a real image. Errors introduced at each step of the inversion process can accumulate, leading to a suboptimal reconstruction. Current methodologies, which focus on optimizing the transition between successive steps, may not be adequate to address this issue holistically. To tackle this, we propose a novel approach that considers the inversion process as a whole, underscoring the significance of the initial latent state throughout the process. Our approach, named EasyInv, incorporates a straightforward mechanism to periodically reinforce the influence of the initial latent state during the inversion. This is realized by blending the current latent state with the previous one at strategically selected intervals, thereby increasing the weight of the initial latent state and diminishing the noise’s impact. As a result, EasyInv ensures a reconstructed version that remains closer to the original image, as illustrated in Figure 1(d). Furthermore, by building upon the traditional DDIM Inversion framework (Couairon et al., 2023), EasyInv does not depend on iterative optimization between adjacent steps, thus enhancing computational efficiency. In Figure 2, we present a visualization of the latent states at the midpoint of the total denoising steps for various inversion methods. The outcomes of our EasyInv are more closely aligned with the original image compared to all other methods, demonstrating that EasyInv achieves faster convergence.
@@
Figure 2: Visualization of latent states at midway denoising steps for various inversion methods. Our EasyInv shows enhanced convergence by closely approximating the original image.
2 Related Works
@@
Diffusion Model. In recent years, there has been significant progress in the field of generative models, with diffusion models emerging as a particularly popular approach. The seminal denoising diffusion probabilistic models (DDPM) (Ho et al., 2020) introduced a practical framework for image generation based on the diffusion process. This method stands out from its predecessors, such as generative adversarial networks (GANs), due to its iterative nature. During the data preparation phase, Gaussian noise is incrementally added to a real image until it transitions into a state that is indistinguishable from raw Gaussian noise. Subsequently, a model can be trained to predict the noise added at each step, enabling users to input any Gaussian noise and obtain a high-quality image as a result. Ho et al(Ho et al., 2020) provided a robust theoretical foundation for their model, which has facilitated further advancements. Generative process in DDPM is both time-consuming and inherently stochastic due to the random noise introduced at each step. To address these limitations, the denoising diffusion implicit models (DDIM) were developed (Song et al., 2020). By reformulating DDPM, DDIM has reduced the amount of random noise added at each step. This reformulation results in a more deterministic denoising process. Furthermore, the absence of random noise allows for the aggregation of several denoising steps, thereby reducing the overall computation time required to generate an image.
@@
Image Inversion. Converting a real image into noise is pivotal in real image editing using diffusion models. The precision of this process has a profound impact in the final edit, with the critical element being the accurate identification of the noise added at each step. Couairon et al(Couairon et al., 2023) swapped the roles of independent and implicit variables within the denoising function of the DDIM model, enabling it to predict the noise that should be introduced to the current latents. However, it is essential to recognize that the denoising step in a diffusion model is inherently an approximation, and when this approximation is utilized inversely, discrepancies between the model’s output and the actual noise value are likely to be exacerbated. To address this issue, ReNoise (Garibi et al., 2024) iterates through each noising step multiple times. For each inversion step, they employ an iterative approach to add and subsequently reduce noise, with the noise reduced in the final iteration being carried forward to the subsequent iteration. Pan et al. (Pan et al., 2023) offered a theoretical underpinning to the ReNoise. Iterative optimization from ReNoise is classified under the umbrella of fixed-point iteration methods. Building upon Anderson’s seminal work (Anderson, 1965), Pan et al. have advanced the field by proposing their novel method for optimizing noise during the inversion process.
3 Methodology
3.1 Preliminaries
@@
DDIM Inversion. Let $\mathbf{z}_{T}$ denote a noise tensor with $\mathbf{z}_{T}\sim\mathcal{I}(0,\mathbf{I})$. The DDIM (Couairon et al., 2023) leverages a pre-trained neural network $\varepsilon_{\theta}$ to perform $T$ denoising diffusion steps. Each step aims to estimate the underlying noise and subsequently restore a less noisy version of the tensor, $\mathbf{z}_{t-1}$, from its noisy counterpart $\mathbf{z}_{t}$ as:
@@
$\begin{split}\mathbf{z}_{t-1}&=\sqrt{\frac{\alpha_{t-1}}{\alpha_{t}}}\mathbf{z}_{t}+\\ &\Bigg{(}\sqrt{\frac{1}{\alpha_{t-1}}-1}-\sqrt{\frac{1}{\alpha_{t}}-1}\Bigg{)}\cdot\varepsilon_{\theta}\big{(}\mathbf{z}_{t},t,\tau_{\theta}(y)\big{)},\end{split}$ (1)
@@
where $t=T\rightarrow 1$, and $\{\alpha_{t}\}_{t=1}^{T}$ constitutes a prescribed variance set that guides the diffusion process. Furthermore, $\tau_{\theta}$ serves as an intermediate representation that encapsulates the textual condition $y$. We denote:
@@
Re-evaluating Equation 1, we derive DDIM Inversion process (Couairon et al., 2023) as presented in Equation 3. In this reformulation, we relocate an approximate $\mathbf{z}^{*}_{t}$ to the left-hand side, resulting in the following expression:
@@
Review. Given an image $\mathbf{I}^{*}$, after encoding it into the latent $\mathbf{z}^{*}_{0}$, we initiate $T$ inversion steps using Equation 3 to obtain the noise $\mathbf{z}^{*}_{T}$. Starting with $\mathbf{z}_{T}=\mathbf{z}^{*}_{T}$, we proceed with a denoising process in Equation 1 to infer an approximate reconstruction $\mathbf{z}_{0}$ that resembles the original latent $\mathbf{z}^{*}_{0}$. The primary source of error in this reconstruction arises from the difference between the noise predicted during the inversion process $\varepsilon_{\theta}\big{(}\mathbf{z}^{*}_{t-1},t-1,\tau_{\theta}(y)\big{)}$ and the noise expected in the sampling process, $\varepsilon_{\theta}\big{(}\mathbf{z}_{t},t,\tau_{\theta}(y)\big{)}$, denoted as $\varepsilon_{t}$, at each iterative step. This discrepancy originates from an imprecise approximation of the time step from $t$ to $t-1$. Therefore, reducing the discrepancy between the predicted noises at each step is crucial for achieving an accurate reconstruction, which is essential for the success of subsequent image editing tasks. For simplicity below, we define:
$\varepsilon_{t}^{*}=\varepsilon_{\theta}\big{(}\mathbf{z}^{*}_{t-1},t-1,\tau_{\theta}(y)\big{)},\;\varepsilon_{t}=\varepsilon_{\theta}\big{(}\mathbf{z}_{t},t,\tau_{\theta}(y)\big{)}.$ (4)
3.2 Fixed-Point Iteration
@@
For clarity, let’s first restate Equation 3 as follows:
@@
which represents the introduction of adding noise to the latent state $\mathbf{z}^{*}_{t-1}$. Under the assumption of Equation 5, it should be the case that:
@@
Subsequently, by employing the noise estimation function from Equation 2, we obtain:
@@
Combining $d(\mathbf{z}_{t})=\varepsilon_{t}$ and Equation 5, we obtain:
@@
This formulation presents a fixed-point problem, which pertains to a value that remains unchanged under a specific transformation (Bauschke et al., 2011). In the context of functions, a fixed point is an element that is invariant under the application of the function. We seek a $\varepsilon^{*}_{t}$ that, when transformed by $g$ and followed by $d$, can map back to itself, signifying an optimal solution as per Equation 5.
@@
where $n$ denotes the iteration count. This iterative process can be enhanced through acceleration techniques such as Anderson acceleration  (Anderson, 1965). However, calculating a complex $\varepsilon^{*}_{t}$ can be quite onerous. An empirical acceleration method proposed (Pan et al., 2023) introduces a refinement for $\varepsilon_{t}^{*}$ by setting: $\big{(}\varepsilon^{*}_{t})^{n}=\varepsilon_{\theta}(({\mathbf{z}^{*}_{t}})^{n},t-1,\tau_{\theta}(y)\big{)}$ and $(\varepsilon^{*}_{t})^{n-1}=\varepsilon_{\theta}\big{(}({\mathbf{z}^{*}_{t}})^{n-1},t-1,\tau_{\theta}(y)\big{)}$. They finally reach:
@@
where the term $0.5\cdot(\varepsilon^{*}_{t})^{n}+0.5\cdot(\varepsilon^{*}_{t})^{n-1}$ represents the refinement technique for $\varepsilon^{*}_{t}$ as suggested by Pan et al. If we were to apply the function $d$ to both sides of Equation 11, it would align perfectly with the form of Equation 10. Their experiments have demonstrated that this approach is more effective than both Anderson’s method  (Anderson, 1965) and other techniques in inversion tasks.
Despite the progress made, this paper acknowledges inherent limitations in the practical implementation of the inversion technique. (1) Inversion Efficiency: While the method outlined in Equation 11 has shown improvements over traditional fixed-point iteration, it still relies on iterative optimization. The need for multiple forward passes through the diffusion model is computationally demanding and can result in inefficiencies in downstream applications. (2) Inversion Performance: The theoretical improvements presented assume that $\varepsilon^{*}_{t}=\varepsilon_{t}$. However, iterative optimization does not guarantee the exact fulfillment of Equation 7 for every time step $t$. Therefore, while the method may theoretically offer superior performance, cumulative errors can sometimes lead to practical outcomes that are less satisfactory than those achieved with the standard DDIM Inversion method, as shown in Figure 1.
3.3 Kalman Filter
The Kalman filter is designed to estimate the state of a dynamic system from noisy measurements. It assumes two states of the system: the predicted state, $x_{k}$, which is calculated using the process function, and the observed state, $y_{k}$, which represents the measured state of the system. For instance, the position of an object could be predicted using its velocity and time, while its position might be observed via a radar. Here, $k$ denotes the time-step. Ideally, $x_{k}$ and $y_{k}$ should be identical; however, in practice, both are prone to inaccuracies. The Kalman filter aims to provide an optimized estimate by combining both values. The relevant equations are as follows:
$x_{k}=Ax_{k-1}+Bu_{k}+w_{k-1}.$ (12)
$y_{k}=Hx_{k}+v_{k}.$ (13)
In these equations, $x_{k}$ represents the true value at time-step $k$, which we aim to estimate; $u_{k-1}$ is the control input at time step $k-1$, such as acceleration; $w_{k-1}$ and $v_{k}$ are random noise or errors introduced during calculation and measurement, thus $y_{k}$ represents the measurement value; $A$ and $B$ are weight matrices; and $H$ is the transformation matrix. When $x_{k}$ and $y_{k}$ are of the same type, the matrix $H$ can be omitted. In practice the value of $w_{k-1}$ is unknown, otherwise the precise results would be calculated without the needs of Kalman filter. Instead, we have the predicted value $\bar{x}_{k}$, which is $\bar{x}_{k}=Ax_{k-1}+Bu_{k}$. The key idea of the Kalman filter is to fuse these $\bar{x}_{k}$ and $y_{k}$ in order to obtain a more accurate estimation of the system state. The fusion process is typically represented as:
$\tilde{x}_{k}=\alpha\bar{x}_{k}+\beta{y^{measure}},$ (14)
where $\alpha+\beta=I$ and $y^{measure}=H^{-1}y_{k}$. We can rewrite the fusion equation as:
$\tilde{x}_{k}=(1-\beta)\bar{x}_{k}+\beta{H^{-1}y_{k}}=\bar{x}_{k}+\beta(H^{-1}y_{k}-\bar{x}_{k}),$ (15)
When we set $\beta=K\cdot H$ the objective function would be:
$\tilde{x}_{k}=\bar{x}_{k}+K(y_{k}-H\bar{x}_{k}).$ (16)
Here, $K$ is the Kalman gain. Additionally, $\bar{x}_{k}=A\tilde{x}_{k-1}+Bu_{k}$ is the predicted estimate of the true state $x_{k}$, based on the previous estimate $\tilde{x}_{k-1}$.
3.4 EasyInv
To facilitate our subsequent analysis, we introduce the notation $\bar{\alpha}_{t}$ to represent $\sqrt{\frac{\alpha_{t}}{\alpha_{t-1}}}$ and $\bar{\beta}_{t}$ to denote $\sqrt{\frac{\alpha_{t}}{\alpha_{t-1}}}\Big{(}\sqrt{\frac{1}{\alpha_{t-1}}-1}-\sqrt{\frac{1}{\alpha_{t}}-1}\Big{)}$. With these notations, we can reframe Equation 3 as follow:
$\mathbf{z}^{*}_{t}=\bar{\alpha}_{t}\mathbf{z}^{*}_{t-1}+\bar{\beta}_{t}\varepsilon^{*}_{t}.$ (17)